OpenStudy (anonymous):

Need help! :( The sum of 4 integers in A.P is 24, and their product is 945. Find the 4 integers.

OpenStudy (anonymous):

since it is in AP it will be wise to take the 4 integers in AP as a-3d,a-d,a+d,a+3d now by the question

OpenStudy (anonymous):

a-3d+a-d+a+d+a+3d=24 -->> 4a=24 --> a=6 also (a-3d)(a-d)(a+d)(a+3d)=495 or (a^2-9d^2)(a^2-d^2)=495 by putting the value of a=6 in the above eq we will be able to get d thus if a and d are known th e reqd integers are known

OpenStudy (whpalmer4):

we know that one of the numbers must be a multiple of 5 to have the product be divisible by 5 4+5+6+7 = 22, 5+6+7+8 = 26, so need to spread them out but with bias to making sum larger. try 3 + 5 + 7 + 9 = 24

OpenStudy (whpalmer4):

matricked reversed two of the digits, so it should be (a-3d)(a-d)(a+d)(a+3d)=945 if you are going to go that route

OpenStudy (anonymous):

uhm what is 'A.P'

OpenStudy (whpalmer4):

A.P. = arithmetic progression